OpenStudy (anonymous):

f(x)=x^2+6x-16, I need to convert this to vertex form. HELP PLS!!!!!!

OpenStudy (zale101):

do you know how to complete the square?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

I was looking on youtube for explanations but I couldnt find any

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

Is the bottom the answer?

OpenStudy (anonymous):

cause I need to explain how to solve it

OpenStudy (zale101):

\[x^2+6x-16=y\] \[x^2+6x-16+16=y+16\] \[(x^2+6x)=y+16\]Take half of the coefficient and then square, the coefficient in this case is 6

OpenStudy (zale101):

can you ake half of the coeffeincet and then square it? \((\frac{ 6 }{ 2 })^2\)

OpenStudy (zale101):

take*

OpenStudy (zale101):

\[(x^2+6x+(\frac{ 6 }{ 2 })^{2})=y+16\]

OpenStudy (anonymous):

Do I solve this and get the answer? @Zale101

OpenStudy (zale101):

That's just part of the process.

OpenStudy (anonymous):

dang

OpenStudy (zale101):

(6/2)^2=?